OpenStudy (anonymous):

I need help with integral of 2x - 3cos2x please

OpenStudy (anonymous):

specifically the trig...

OpenStudy (trojanpoem):

Notice that the Integration of cosx = sinx, x^n = x^n+1/ n+1 , Let's start : \[\int\limits_{}^{} 2x - 2\cos2x = \frac{ 2x^2 }{ 2 } - \frac{ 2\sin(2x) }{ 2 } + c\] \[x^2 - \sin(2x) + c\]

OpenStudy (trojanpoem):

After integrating , it's recommended to check if your answer is right ( 99.5 this check works) get the derivative. If both the question and your derivative are equal be sure you are.

OpenStudy (anonymous):

looks like you may have misread the problem? Here's what wolfram gave us: http://www.wolframalpha.com/input/?i=integral+of+2x+-+3cos2x

OpenStudy (trojanpoem):

my bad just replace 2/2 with 3/2 so you have x^2 - 3/2 sin2x + c wolframalpha just used double angle rule sin2x = 2sinxcos so x^2 - 3sinxcosx + c

OpenStudy (trojanpoem):

To me , keep it sin(2x)

OpenStudy (anonymous):

@TrojanPoem , can you break it down further?

OpenStudy (welshfella):

No you cant break it down any further

OpenStudy (anonymous):

got it, beginning to reason through the heuristics of inverse reasoning, although slowly